A particle is placed inside a one
dimensional box with rigid walls. The first
excited state energy of the particle is
4 eV. The ground state energy is
(A) 2 eV (B) 3 eV
(C) 1 eV (D) 0 eV
1
Expert's answer
2019-04-08T12:56:21-0400

(C) 1 eV




"\frac {W^2_1} {W^2_0}=\frac {2^2} {1^2} (1)"

Using (1) we got: W0=1